OpenStudy (anonymous):

Mary and Laci go to the Mexican restaurant. Mary orders 2 tacos and 2 enchiladas. Laci orders 3 tacos and 1 enchilada. Mary’s bill was $4.80. Laci’s bill was $4.00. What was the price of the taco. $0.80 $1.60 $2.40 $3.00

OpenStudy (anonymous):

2T + 2E = 4.8 3T + 1E = 4 multiple ( -2 ) ......................... 2T + 2E = 4.8 -6T - 2E = -8 adding -4T = -3.2 4T = 3.2 T = 0.8

OpenStudy (anonymous):

thanks love (: